What is the change in kinetic energy (in units of MJ) associated with an aircraft of mass 353 metric tonnes changing its speed from 121 miles-per-hour to 369 miles-per-hour?
Answer #1

u = 121 mph = 54.1 m/s

v = 369 mph = 164. 96 m/s

m = 3.53*10^5 kg

chabge in kE

dkE = 0.5 m ( v^2 - u^2)

dKE = 0.5* 3.53*10^5* ( 164.96^2 - 54.1^2)

dKE = 4.286* 10^3 MJ

======

Comment in case any doubt, will reply for sure.. Goodluck